Calculate the number density of valence electrons in lithium given that the mass density is 530kg/m the molar mass is 7g/mol and that lithium is monovalent assuming that the mean free path of the electron in lithium is 1 nanometer calculate the resistivity of lithium at 295k
1
Expert's answer
2019-08-27T09:34:05-0400

The number density of valence electrons is given by formula

"n=\frac{N_A Zρ}{M} (1)"

where NA is the Avogadro’s constant, Z is the valency, ρ is the density, M is the the molar mass

Using (1) we got

In our case, we have NA=6.02×1023 mol-1, Z=1, ρ=530kg/m3, M=7×10-3 kg/mol

n= 4.54 x 1028 m-3

Answer

4.54 x 1028 m-3



The resistivity is is given by formula

"ρ=\frac{m}{q×q×n×τ} (2)"

where m is the effective mass of the conduction electrons (m=9.1×10-31 kg), q is the elementary charge (q=1.6×10-19 C), n is the density of conduction electrons (n=4.54 x 1028 m-3), τ is the mean time between collisions.


Mean free path is given by

"d=\u03c4\u00d7v (3)"

where v is the Fermi velocity for Lithium


The Fermi velocity is given by formula

"v= \sqrt\frac{3kT} {m} (4)"

where k is the Boltzmann constant, m is the effective mass of the conduction electrons (m=9.1×10-31 kg), T=295 K


Using (4) we got

v=116×105 m/s (5)


We can find the mean time between collisions

τ=8.6×10-17 s (6)


We put (6) in (2)

ρ=9.4×10-8 m×Ώ


Answer

9.4×10-8 m×Ώ